OpenStudy (loser66):

Suppose h(x) and g(x) differentiable at x =a Let f(x) = g(x) if x>=a f(x) = h(x) is x

OpenStudy (loser66):

\(f(x) =\begin{cases}g(x) ~~if~~x\geq a\\h(x)~~if~~x<a\end{cases}\)

OpenStudy (misty1212):

HI!

OpenStudy (loser66):

Hello, help me, please

OpenStudy (misty1212):

not sure what there is to show you need to have \(f\) be continuous, which is the same as saying \[g(a)=h(a)\]

OpenStudy (loser66):

Yes, it is

OpenStudy (misty1212):

i mean that is one part you can drag it out \[\lim_{x\to a^+}f(x)=g(a)\\ \lim_{x\to a ^-}f(x)=g(a)\] so if \(g(a)=h(a)\) then the limit exists and is that number

OpenStudy (misty1212):

i made a mistake on the second line it should be \(h(a)\)

OpenStudy (loser66):

Yes, how about the opposite direction?

OpenStudy (misty1212):

guess you are not finished though you also have to say the derivatives are the same too right?

OpenStudy (misty1212):

oh but you are given that they are, so there is not too much to say ...

OpenStudy (loser66):

I think derivative part is for the opposite direction, right?

OpenStudy (loser66):

oh, I got it

OpenStudy (loser66):

Thanks for the help

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]